OpenStudy (anonymous):

solve the following using addition... 2/3 x- 4/3y = 1/2 -4x-8y=10

OpenStudy (anonymous):

did u try to do this problem

OpenStudy (anonymous):

struggling with this section

OpenStudy (anonymous):

Angela did one for you.... I'll give u a hint : multiply the first equation by -6 and then add second to the result

OpenStudy (anonymous):

2/3 x- 4/3y = 1/2 -------------- (1) -4x-8y=10 ---------------- (2) multiply (1) by -6 and tell me what u get...

OpenStudy (anonymous):

-4x+8y=-3

OpenStudy (anonymous):

so y= -7/8

OpenStudy (anonymous):

yes then solve it rhoff

OpenStudy (anonymous):

No, no, how can u get y = -7/8????

OpenStudy (anonymous):

call the equation u got as (3) now add (3) and (2), n see what u get

OpenStudy (anonymous):

-4x-8y=10 ------------- (2) -4x+8y=-3 --------------(3)

OpenStudy (anonymous):

-8x=7

OpenStudy (anonymous):

x=-7/8

OpenStudy (anonymous):

Correct !!! so put x = -7/8 in (2) to get value of y...

OpenStudy (anonymous):

i get y=-27/8

OpenStudy (anonymous):

check again...

OpenStudy (anonymous):

sry y= 27/8

OpenStudy (anonymous):

-4x-8y=10 ------------- (2) -7 -4 * ----- - 8y = 10 8 28 7 -- - 8y = 10 ==> ---- - 8y = 10 8 2 7 20 - 7 13 -8y = 10 - ----- ==> -8y = --------- ==> -8y = ----- 2 2 2 13 13 - 13 y = -------- = ------ = ----- 2 * -8 -16 16

OpenStudy (anonymous):

is it clear now????

OpenStudy (anonymous):

it is - and then i express the answer as (-7/8, -13/16) - or do I need to make them whole numbers?

OpenStudy (anonymous):

Yes, u put the answer in (x,y) form as (-7/8, -13/16)

OpenStudy (anonymous):

thank you so much. for some reason anytime fractions are included I really struggle. All the examples the teacher (and the book, and tutor sites) use are nice and clean - it's hard to see if I'm doing it right when the examples don't reflect the actual problems.

OpenStudy (anonymous):

I guess now u hv learnt how to handle fractions !!!! :D Don't come here asking for answers. State clearly that u need to be guided how to solve the question.....